# Physics-Guided Learning of Meteorological Dynamics for Weather Downscaling and Forecasting

### v1

@@ Abstract.
Physics, Learning, Weather, Prediction, Parameterization
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaccs: Applied computing Physical sciences and engineeringccs: Computing methodologies Machine learning
@@ 1. Introduction
Weather prediction remains one of modern science’s most complex and vital challenges, with the nonlinear interactions between various meteorological variables, the vast spatial and temporal scales involved, and the chaotic nature of weather systems (Holton, 1973). The first-principle approach to weather prediction, i.e., numerical weather prediction (NWP), relies on mathematical models of atmospheric and oceanic phenomena. NWP is often computationally intensive at high resolution, and many unresolved physical processes such as precipitation and radiation need to be represented by parameterization. Therefore, there has been a growing interest in using machine learning (ML) models for weather prediction. Deep learning models, trained by nearly 40 years of European Center for Medium-Range Weather Forecasts (ECMWF) reanalysis v5 (ERA5) data (Hersbach et al., 2020), have demonstrated remarkable ability to capture complex nonlinear relationships for tasks such as weather forecasting (Pathak et al., 2022; Wu et al., 2023) and downscaling (Vandal et al., 2017). However, despite the recent success of ML techniques, the application of deep learning to weather prediction is not without challenges. Existing ML models do not incorporate established physical laws (e.g., fluid dynamics, thermodynamics) to ensure that the derived variables are consistent with these laws.
@@
In light of this challenge, we propose a novel framework, PhyDL-NWP. This proposed paradigm first trains a neural network to predict weather conditions based on spatio-temporal coordinates, then leverages automatic differentiation for calculating partial differential equation (PDE) terms. Based on this, we construct a library of physics terms derived from first-principle equations that can be represented using available dataset variables. To address unresolved processes and missing variables, we adopt a parametrization strategy (Warner, 2010), introducing a latent force model as a parametrization term to capture the effects of physical forces not explicitly represented in the constructed PDEs.
@@ 2.1. Weather Prediction
In the area of weather prediction (Alley et al., 2019), Numerical Weather Prediction (NWP) (Lorenc, 1986; Bauer et al., 2015) is the current mainstream method. It uses mathematical models of the atmosphere and oceans, such as partial differential equations (PDE), to predict future weather based on current weather conditions. Some notable NWP models include European Centre for Medium-Range Weather Forecasts (ECMWF) [^1] https://www.ecmwf.int/, Global Forecast System (GFS) [^2] https://www.ncei.noaa.gov/products/weather-climate-models/global-forecast, etc. NWP can forecast weather in the medium range (beyond a few days ahead) but usually involves extensive computation. For example, ECMWF operates one of the largest supercomputer complexes in Europe.
@@
Recently, deep learning has emerged as another promising solution to weather forecasting (Hu et al., 2021; Price et al., 2023) and downscaling (Vandal et al., 2017; Park et al., 2022) tasks, owing to its incredible ability to model complex nonlinear relationships. These deep learning models (Wang et al., 2019; Han et al., 2022) mainly rely on different neural architectures such as LSTM (Li et al., 2022), CNN (Weyn et al., 2020), GNN (Lin et al., 2022; Keisler, 2022) and Transformer (Wu et al., 2023) to capture the evolving dynamics and correlation across space and time. Many large models have emerged in recent years. For example, ClimaX (Nguyen et al., 2023a), GraphCastNet (Lam et al., 2022), ClimateLearn (Nguyen et al., 2023b), FengWu (Chen et al., 2023), Pangu-Weather (Bi et al., 2023) all use backbones such as the Vision Transformer (ViT) (Dosovitskiy et al., 2020), UNet (Ronneberger et al., 2015) and autoencoders, for training a large model for weather forecasting. WeatherBench (Rasp et al., 2020, 2023) benchmarks the use of pre-training techniques for weather forecasting. In addition, FourcastNet (Pathak et al., 2022) leverages the adaptive Fourier neural operator (AFNO) (Li et al., 2020; Guibas et al., 2021) to treat weather as a latent PDE system. There are also works on the prediction of specific variables, such as precipitation (Gao et al., 2024; Yu et al., 2024).
In recent years, a few studies such as NeuralGCM (Kochkov et al., 2024) and WeatherGFT (Xu et al., 2024) have started to explore the integration of underlying physical mechanisms (Anonymous, 2024) in weather prediction. However, they act as a standalone model rather than a universally adaptive module, and there are still no explicit efforts to use parameterization to ensure the completeness of the primitive equations based on the data.
2.2. Spatio-temporal modeling
Spatio-temporal modeling based on deep learning (Liang et al., 2025; Yu et al., 2021b) has thrived in recent years, showing promising results in various applications such as video recognition (Cai et al., 2021), traffic flow forecasting (Wen et al., 2023a; Luo et al., 2019), disease spread modeling (Arenas et al., 2020), etc. Weather prediction shares striking similarities with these applications. Thus, many previous works (Moosavi et al., 2019; Castro et al., 2021; Han et al., 2021) also approach the weather prediction task from the perspective of spatio-temporal modeling. Many models such as recurrent neural networks (Liu et al., 2016; Wang et al., 2022), convolutional neural networks (Tran et al., 2015; Xu et al., 2019), graph networks (Yu et al., 2018; Geng et al., 2019; Jin et al., 2024), transformers (Law and Lucas, 2022; Wen et al., 2023b) and hybrid models (Yan et al., 2021) are extensively applied to capture the correlations of variables across space and time.
2.3. Dynamical system modeling
@@
Dynamical systems modeling (Morton et al., 2018; Long et al., 2018; Li et al., 2019) involves the mathematical formulation of systems whose states evolve over time, which often takes the form of differential equations for continuous systems. Given the governing equations, physics-informed approaches (Raissi et al., 2019; Karniadakis et al., 2021) use the physics mechanism to enhance the dynamical systems. In the absence of governing equations, the identification of physical equations (Brunton et al., 2016; Seo et al., 2019; Luo et al., 2022, 2023) is proposed to explain the observation data and provide theoretical insights with respect to laws of physics (Wu and Tegmark, 2019; Iten et al., 2020). Symbolic learning (Schmidt and Lipson, 2009; Chen et al., 2021, 2022) and sparse regression (Rudy et al., 2017; Rao et al., 2021) have been extensively studied.
@@ 4.1. Downscaling Performance Comparison
We evaluate the effectiveness of PhyDL-NWP and other baseline models for weather downscaling on a real-world dataset Huadong, which is derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) operational forecast (HRES) and reanalysis (ERA5) archive. It comprises a grid of $64\times 44$ cells, with each cell having a grid size of 0.25 degrees in both latitude and longitude. More data details can be found in Appendix A.1. Since most previous studies on weather downscaling can only handle the downscaling of the two spatial dimensions, for the sake of comparison, we also only report the performance of PhyDL-NWP on spatial downscaling in Table 2. We perform 2x and 4x downscaling tasks with 0.5 and 1 degrees resolutions, respectively. To facilitate this, the 0.25-degree HRES data undergoes linear interpolation to generate the requisite 0.5-degree and 1-degree input data. We compare our model against the Bicubic interpolation (Keys, 1981), FSRCNN (Passarella et al., 2022), ResDeepD (Sharma and Mitra, 2022), EDSR (Jiang and Chen, 2022), RCAN (Yu et al., 2021a), YNet (Liu et al., 2020), DeepSD (Vandal et al., 2017), and GINE (Park et al., 2022). For the deep learning baselines, channel-wise normalization is performed for efficiency. Details about baselines can be found in Appendix A.2.4.
Unlike previous methods that take in coarse-granular data and produce fine-granular super-resolution data, PhyDL-NWP, although trained on coarse-granular data $\mathbb{X}$ to learn to align with physics, does not require fine-granular labels $\mathbb{Y}$. Therefore, while baseline models are trained by the input-label pairs of the training dataset and evaluated by the labels of the test dataset, we train PhyDL-NWP using the test inputs without labels, and evaluate it on the test labels. We consider this setting to be fair because the objective is only to perform super-resolution "imaging", in which it does not matter whether a model is online learning like PhyDL-NWP or learning from historical data like most baselines.
@@ 4.2. Forecasting Performance Comparison
There are five kinds of baseline models in comparison, including: (1) Physics-based models: NWP, PINN (Raissi et al., 2019), PINO (Li et al., 2024); (2) Meteorological models: Bi-LSTM-T (Yang et al., 2022), Hybrid-CBA (Han et al., 2022); (3) Vision models: ConvLSTM (Shi et al., 2015), FourcastNet (Pathak et al., 2022) based on AFNO (Guibas et al., 2021), ClimaX (Nguyen et al., 2023a) based on ViT (Dosovitskiy et al., 2020); (4) Spatio-temporal graph models: MTGNN (Wu et al., 2020), MegaCRN (Jiang et al., 2023), GraphCast (Lam et al., 2022). Some baseline models are slightly modified to adapt to the multi-step prediction setting and/or specific modeling resolution, with details described in the Appendix A.2.4. Besides deep learning models, we also compare these models with the Numerical Weather Prediction (NWP) results provided by ECMWF IFS and the Physical-Informed Neural Network (PINN) (Raissi et al., 2019) based on the PDEs learned by PhyDL-NWP. We denote the baseline models as BaseModels and incorporate them with PhyDL-NWP as BaseModels+.
@@ 4.2.1. Regional Weather Forecasting
To understand the holistic properties of PhyDL-NWP, we conduct detailed analyses on the Ningxia dataset. First, the comparison of different models for different forecasting ranges is visualized in Fig. 3. The comparisons on more weather variables are shown in Figs. 6-9 in Appendix 10. BaseModels+ excels BaseModels and NWP at all time steps. As the forecasting range increases, the deep learning performance decreases quickly. The improvement provided by PhyDL-NWP, however, is increasing in the forecasting range, which highlights its unique advantages of guiding models for long-term medium-range forecasting.
@@ 4.3. Meteorology Dynamics Interpretation
Table 5. Comparison of learned dynamics against ground truth in literature (Holton, 1973). Orange color marks terms unavailable from the data, which are represented by latent force $Q$.
@@
Temperature Learned vs. True PDE Learned $\frac{\partial T}{\partial t}=-U_{10}\frac{\partial T}{\partial x}-V_{10}\frac{\partial T}{\partial y}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial T}{\partial t}=-U\frac{\partial T}{\partial x}-V\frac{\partial T}{\partial y}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}W\frac{\partial T}{\partial z}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}k\frac{\partial^{2}T}{\partial z^{2}}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}H}$ Wind Velocity Learned vs. True PDE Learned $\frac{\partial U_{10}}{\partial t}=-U_{10}\frac{\partial U_{10}}{\partial x}-V_{10}\frac{\partial U_{10}}{\partial y}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial U}{\partial t}=-U\frac{\partial U}{\partial x}-V\frac{\partial U}{\partial y}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}W\frac{\partial U}{\partial z}}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}\frac{1}{\rho}\frac{\partial p}{\partial x}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}\nu\Delta\textbf{u}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}F_{f_{x}}}$ Surface Pressure Learned vs. True PDE Learned $\frac{\partial^{2}p}{\partial t^{2}}=\frac{\partial^{2}p}{\partial x^{2}}+\frac{\partial^{2}p}{\partial y^{2}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial^{2}p}{\partial t^{2}}=\frac{\partial^{2}p}{\partial x^{2}}+\frac{\partial^{2}p}{\partial y^{2}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}\frac{\partial^{2}p}{\partial z^{2}}}$ Humidity Learned vs. True PDE Learned $\frac{\partial q}{\partial t}=-U\frac{\partial q}{\partial x}+\frac{\partial^{2}p}{\partial y^{2}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial q}{\partial t}=-U\frac{\partial q}{\partial x}-V\frac{\partial q}{\partial y}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}W\frac{\partial q}{\partial z}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}\nabla(K_{q}\nabla q)}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}S_{q}}$
@@
To understand how PhyDL-NWP is grounded on laws of physics, we compare the learned dynamics by PhyDL-NWP from Huadong dataset with the basic equations of NWP (Holton, 1973; Stull, 1988; Pu and Kalnay, 2019). These equations originate from conservation of mass, energy, and momentum. PhyDL-NWP will explicitly use the terms that appear in the first-principle equations and can be represented based on the dataset, while modeling the rest of the dynamics using latent force parameterization. As shown in Table 5, our latent parameterization strategy is proposed considering that many terms cannot be modeled explicitly, due to missing records in the data or difficulty to measure.
@@
For the modeling of wind velocity, the compared basic equation originates from the 3D Navier-Stokes equations. The velocity is influenced by convective acceleration, the gradient of pressure $p$ with fluid density $\rho$, the diffusion of momentum due to viscosity $\nu$ for velocity of all directions u, and external force $F_{f_{x}}$ along the $x$ direction (such as friction). It is obvious that the WeatherBench dataset does not provide all the necessary variables to complete the equation. Our latent force model will address this problem well.
Similarly, for the modeling of surface pressure, the learned equation also aligns with the 3D wave equation while the missing vertical component is for the latent force $Q$ to capture. For humidity, the equation originates from the conservation of mass, and combines the advection, diffusion and source terms, where $q$ is the humidity, $K_{q}$ is the diffusion coefficient, and $S_{q}$ is the source term for humidity (e.g., evaporation, condensation).

### v2

@@ Abstract.
Physics; Learning; Weather; Prediction; Parameterization
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737081isbn: 979-8-4007-1454-2/2025/08ccs: Applied computing Physical sciences and engineeringccs: Computing methodologies Machine learning
@@ 1. Introduction
Weather prediction remains one of modern science’s most complex and vital challenges, with the nonlinear interactions between various meteorological variables, the vast spatial and temporal scales involved, and the chaotic nature of weather systems. The first-principle approach to weather prediction, i.e., numerical weather prediction (NWP), relies on mathematical models of atmospheric and oceanic phenomena. NWP is often computationally intensive at high resolution, and many unresolved physical processes such as precipitation and radiation need to be represented by parameterization. Therefore, there has been a growing interest in using machine learning (ML) models for weather prediction. Deep learning models, trained by nearly 40 years of European Center for Medium-Range Weather Forecasts (ECMWF) reanalysis v5 (ERA5) data (Hersbach et al., 2020), have demonstrated remarkable ability to capture complex nonlinear relationships for tasks such as weather forecasting (Pathak et al., 2022; Wu et al., 2023) and downscaling (Vandal et al., 2017). However, despite the recent success of ML techniques, the application of deep learning to weather prediction is not without challenges. Existing ML models do not incorporate established physical laws (e.g., fluid dynamics, thermodynamics) to ensure that the derived variables are consistent with these laws.
@@
In light of this challenge, we propose a novel framework, PhyDL-NWP. This proposed paradigm first trains a neural network to predict weather conditions based on spatio-temporal coordinates, then leverages automatic differentiation for calculating partial differential equation (PDE) terms. Based on this, we construct a library of physics terms derived from first-principle equations that can be represented using available dataset variables. To address unresolved processes and missing variables, we adopt a parametrization strategy, introducing a latent force model as a parametrization term to capture the effects of physical forces not explicitly represented in the constructed PDEs.
@@ 2. Related Work
In the area of weather prediction, Numerical Weather Prediction (NWP) (Lorenc, 1986; Bauer et al., 2015) is the current mainstream method. It uses mathematical models of the atmosphere and oceans, such as partial differential equations (PDE), to predict future weather based on current weather conditions. Some notable NWP models include European Centre for Medium-Range Weather Forecasts (ECMWF) [^1] https://www.ecmwf.int/, Global Forecast System (GFS) [^2] https://www.ncei.noaa.gov/products/weather-climate-models/global-forecast, etc. NWP can forecast weather in the medium range but usually involves extensive computation. For example, ECMWF operates one of the largest supercomputer complexes in Europe.
@@
Recently, deep learning has emerged as another promising solution to weather forecasting (Hu et al., 2021; Price et al., 2023) and downscaling (Vandal et al., 2017; Park et al., 2022) tasks. These deep learning models (Wang et al., 2019; Han et al., 2022) rely on different neural architectures such as LSTM (Li et al., 2022), CNN (Weyn et al., 2020), GNN (Lin et al., 2022) and Transformer (Wu et al., 2023) to capture the evolving dynamics and correlation across space and time. Many large models have emerged in recent years. For example, ClimaX (Nguyen et al., 2023a), GraphCastNet (Lam et al., 2022), ClimateLearn (Nguyen et al., 2023b), FengWu (Chen et al., 2023), Pangu-Weather (Bi et al., 2023) all use backbones such as the Vision Transformer (ViT), UNet and autoencoders, for training a large model for weather forecasting. WeatherBench (Rasp et al., 2020) benchmarks the use of pre-training techniques for weather forecasting. In addition, FourcastNet (Pathak et al., 2022) leverages the adaptive Fourier neural operator (AFNO) (Li et al., 2020; Guibas et al., 2021) to treat weather as a latent PDE system. In recent years, a few studies such as NeuralGCM (Kochkov et al., 2024) and WeatherGFT (Xu et al., 2024) have started to explore the integration of underlying physical mechanisms (Verma et al., 2023) in weather prediction. However, there are still no explicit efforts to use parameterization to ensure the completeness of the primitive equations based on the data.
@@
In addition, spatio-temporal modeling based on deep learning (Liang et al., 2025) has thrived in recent years. Many previous works also approach the weather prediction task from the perspective of spatio-temporal modeling (Han et al., 2021). Dynamical systems modeling involves the formulation of systems whose states evolve over time. Given the governing equations, physics-informed approaches (Raissi et al., 2019; Karniadakis et al., 2021) use the physics mechanism to enhance the dynamical systems. In the absence of governing equations, the identification of physical equations (Luo et al., 2022; Chen et al., 2022; Luo et al., 2023) is proposed to provide insights with respect to the laws of physics.
@@ 4.1. Downscaling Performance Comparison
We evaluate the effectiveness of PhyDL-NWP and other baseline models for weather downscaling on a real-world dataset Huadong, which is derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) operational forecast (HRES) and reanalysis (ERA5) archive. It comprises a grid of $64\times 44$ cells, with each cell having a grid size of 0.25 degrees in both latitude and longitude. More data details can be found in Appendix A.1. Since most previous studies on weather downscaling can only handle the downscaling of the two spatial dimensions, for the sake of comparison, we also only report the performance of PhyDL-NWP on spatial downscaling in Table 2. We perform 2x and 4x downscaling tasks with 0.5 and 1 degrees resolutions, respectively. To facilitate this, the 0.25-degree HRES data undergoes linear interpolation to generate the requisite 0.5-degree and 1-degree input data. We compare our model against the Bicubic interpolation, FSRCNN (Passarella et al., 2022), ResDeepD (Sharma and Mitra, 2022), EDSR (Jiang and Chen, 2022), RCAN (Yu et al., 2021), YNet (Liu et al., 2020), DeepSD (Vandal et al., 2017), and GINE (Park et al., 2022). For the deep learning baselines, channel-wise normalization is performed for efficiency. Unlike prior methods that rely on fine-granular labels for supervision, PhyDL-NWP learns from coarse-granular inputs alone by aligning with physical principles, enabling it to perform super-resolution without labeled training outputs. Details about baselines can be found in Appendix A.2.4.
@@ 4.2. Forecasting Performance Comparison
There are five kinds of baseline models in comparison, including: (1) Physics-based models: NWP, PINN (Raissi et al., 2019), PINO (Li et al., 2024); (2) Meteorological models: Bi-LSTM-T (Yang et al., 2022), Hybrid-CBA (Han et al., 2022); (3) Vision models: ConvLSTM (Shi et al., 2015), FourcastNet (Pathak et al., 2022) based on AFNO (Guibas et al., 2021), ClimaX (Nguyen et al., 2023a); (4) Spatio-temporal graph models: MTGNN (Wu et al., 2020), MegaCRN (Jiang et al., 2023), GraphCast (Lam et al., 2022). Some baseline models are slightly modified to adapt to the multi-step prediction setting and/or specific modeling resolution, with details described in the Appendix A.2.4. Besides deep learning models, we also compare these models with the Numerical Weather Prediction (NWP) results provided by ECMWF IFS and the Physical-Informed Neural Network (PINN) (Raissi et al., 2019) based on the PDEs learned by PhyDL-NWP. We denote the baseline models as BaseModels and incorporate them with PhyDL-NWP as BaseModels+.
@@ 4.2.1. Regional Weather Forecasting
To understand the holistic properties of PhyDL-NWP, we conduct detailed analyses on the Ningxia dataset. First, the comparison of different models for different forecasting ranges is visualized in Fig. 3. BaseModels+ excels BaseModels and NWP at all time steps. As the forecasting range increases, the deep learning performance decreases quickly. The improvement provided by PhyDL-NWP, however, is increasing in the forecasting range, which highlights its unique advantages of guiding models for long-term forecasting.
@@ 4.3. Meteorology Dynamics Interpretation
Table 6. Comparison of learned dynamics against ground truth in literature. Orange color marks terms unavailable from the data, which are represented by latent force $Q$.
@@
Temperature Learned vs. True PDE Learned $\frac{\partial T}{\partial t}=-U_{10}\frac{\partial T}{\partial x}-V_{10}\frac{\partial T}{\partial y}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial T}{\partial t}=-U\frac{\partial T}{\partial x}-V\frac{\partial T}{\partial y}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}W\frac{\partial T}{\partial z}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}k\frac{\partial^{2}T}{\partial z^{2}}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}H}$ Wind Velocity Learned vs. True PDE Learned $\frac{\partial U_{10}}{\partial t}=-U_{10}\frac{\partial U_{10}}{\partial x}-V_{10}\frac{\partial U_{10}}{\partial y}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial U}{\partial t}=-U\frac{\partial U}{\partial x}-V\frac{\partial U}{\partial y}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}W\frac{\partial U}{\partial z}}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}\frac{1}{\rho}\frac{\partial p}{\partial x}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}\nu\Delta\textbf{u}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}F_{f_{x}}}$ Surface Pressure Learned vs. True PDE Learned $\frac{\partial^{2}p}{\partial t^{2}}=\frac{\partial^{2}p}{\partial x^{2}}+\frac{\partial^{2}p}{\partial y^{2}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial^{2}p}{\partial t^{2}}=\frac{\partial^{2}p}{\partial x^{2}}+\frac{\partial^{2}p}{\partial y^{2}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}\frac{\partial^{2}p}{\partial z^{2}}}$ Humidity Learned vs. True PDE Learned $\frac{\partial q}{\partial t}=-U\frac{\partial q}{\partial x}-V\frac{\partial q}{\partial y}+\frac{\partial^{2}q}{\partial x^{2}}+\frac{\partial^{2}q}{\partial y^{2}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}Q}$ Ground Truth $\frac{\partial q}{\partial t}=-U\frac{\partial q}{\partial x}-V\frac{\partial q}{\partial y}-{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}W\frac{\partial q}{\partial z}}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}k\nabla^{2}q}+{\color[rgb]{0.9,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.55,0}S_{q}}$
@@
To understand how PhyDL-NWP is grounded on laws of physics, we compare the learned dynamics by PhyDL-NWP from Huadong dataset with the basic equations of NWP (Stull, 1988). These equations originate from conservation of mass, energy, and momentum. PhyDL-NWP will explicitly use the terms that appear in the first-principle equations and can be represented based on the dataset, while modeling the rest of the dynamics using latent force parameterization. As shown in Table 6, our latent parameterization strategy is proposed considering that many terms cannot be modeled explicitly, due to missing records in the data or difficulty to measure.
@@
For wind velocity, the compared basic equation originates from the 3D Navier-Stokes equations. The velocity is influenced by convective acceleration, the gradient of pressure $p$ with fluid density $\rho$, the diffusion of momentum due to viscosity $\nu$ for velocity of all directions u, and external force $F_{f_{x}}$ along the $x$ direction (such as friction). It is obvious that the WeatherBench dataset does not provide all the necessary variables to complete the equation. Our latent force model will address this problem well. Similarly, for surface pressure, the learned equation aligns with the 3D wave equation while the missing vertical component is for the latent force $Q$ to capture. For humidity, the equation originates from the conservation of mass, and combines the advection, diffusion and source terms, where $q$ is the humidity, and $S_{q}$ is the source term for humidity (e.g., evaporation, condensation). We approximate precipitation using a parameterized PDE for humidity as a tractable surrogate that captures its dominant drivers.